# A Calculus of Overlays

### v1

@@ 8. Future Work
The Overlay language (Yang, 2025b) is an executable implementation of Overlay-Calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language—spanning both theoretical foundations and library-level encodings.
@@
The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks—proving that well-typed programs do not produce dangling references at runtime—is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how Overlay-Calculus relates to the $\lambda$-calculus: a Typed Overlay-Calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve. In untyped Overlay-Calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary—it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
@@
Overlay-Calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect—it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 4.1). However, many operations (e.g., equality testing) assume that values are linear—inhabiting exactly one constructor. The observer pattern (Appendix B) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
@@
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields. This does not restrict inheritance itself—it remains unconstrained—but provides a way to detect invariant violations before they propagate.
@@
The same technique—observer returning Boolean, followed by Boolean dispatch—yields closed pattern matching. The observer pattern of Appendix B is inherently open: new constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute—they have a fixed order—and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
The trie operations in Appendix A are hardcoded for binary natural number keys. Since every overlay is a trie (Appendix A) and insertion is simply inheritance, only operations that cannot be expressed as inheritance need explicit implementation: deletion, dynamic lookup, and prefix lookup. Generalizing to arbitrary key types requires each key factory to provide two interfaces—SelectTrieChild (select a subtrie by constructor) and ReplaceTrieChild (replace a subtrie by constructor while preserving the rest)—corresponding to the getter and setter of a lens (Foster et al., 2007) focused on the constructor-determined child. Prefix lookup is structurally analogous to equality testing: both synchronously traverse a key (a single-path trie) alongside another structure, dispatching at each level. The difference is the return type—a Boolean for equality, a subtrie for prefix lookup.
Appendix A Trie Operations
@@
$\mathrm{Zero}$ $\;\longrightarrow\;\mathrm{value}$
$\mathrm{Odd}(n)$ $\;\longrightarrow\;\mathrm{at\_odd},\;\text{then path for }n$
$\mathrm{Even}(n)$ $\;\longrightarrow\;\mathrm{at\_even},\;\text{then path for }n$
Write (insertion)
Inserting an entry with value $v$ at key $k$ is inheriting a singleton trie whose nesting mirrors the bit structure of $k$. For example, inserting at key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$:
$\mathrm{trie}^{\prime}=\{\mathrm{trie},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
Each inheritance creates a new immutable trie; the original is unchanged.
Read (static lookup)
When the key is known statically, retrieving the entry at key $k$ is reference-path navigation: the reference follows the same label sequence that was used for insertion. For the same key $5$:
$\mathrm{enclosing\_scope}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$
where $\mathrm{enclosing\_scope}$ is the qualified-this scope name containing the trie definition. The reference mechanism traverses the nested record structure—which is the trie—returning the value at the final node.
Read (dynamic lookup)
When the key is a BinNat value constructed at runtime via $\mathrm{Zero}$, $\mathrm{Odd}$, and $\mathrm{Even}$, the lookup path must be computed dynamically. This is achieved by folding the BinNat key into a Selector—an overlay that, given a trie, navigates to the appropriate node and returns the stored entry. A Selector declares the trie shape and a result slot:
$\mathrm{Selector}\mapsto\{\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;\mathrm{result}\mapsto\{\}\}$
Three observer callbacks build Selectors bottom-up. These callbacks ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) are interpretation points in the sense of Appendix G: they can be filled externally by an orchestrator (CPS interpretation) or inherited locally (direct interpretation), demonstrating the CPS-agnostic property. The $\mathrm{on\_zero}$ callback returns a Selector that reads $\mathrm{value}$ from the trie:
$\mathrm{lookup\_on\_zero}\mapsto\{\mathrm{Selector},\;\mathrm{result}\mapsto\mathrm{lookup\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{value}\}$
The $\mathrm{on\_odd}$ callback takes a Selector (via the $\mathrm{argument}/\mathrm{result}$ encoding of Section 4) and returns a new Selector that navigates to $\mathrm{at\_odd}$, then delegates to the argument Selector:
@@
$\mathrm{lookup\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Selector},\;\mathrm{result}\mapsto\{\mathrm{Selector},\;$
$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;\mathrm{trie}\mapsto\mathrm{lookup\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\}\}$
The $\mathrm{on\_even}$ callback is symmetric, navigating to $\mathrm{at\_even}$ instead.
The $\mathrm{Lookup}$ combinator folds a BinNat key with these callbacks to obtain a Selector, then applies it to the trie:
$\mathrm{Lookup}$ $\mapsto\{$
$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
$\quad\mathrm{trie\_input}\mapsto\{\},$
$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{lookup\_on\_zero},\;$
$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{lookup\_on\_odd},\;$
$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{lookup\_on\_even}\},$
$\quad\mathrm{applied\_selector}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
$\quad\mathrm{result}\mapsto\mathrm{applied\_selector}.\mathrm{result}\}$
For example, looking up key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ folds into a Selector that navigates $\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$—the same path as the static reference, but computed from the key value. The entire construction uses only records (with inheritance sources) and references; no additional primitives are needed.
Delete
Since inheritance is additive—it can only merge properties, never remove them—deletion cannot be expressed as a single inheritance. Instead, deletion reconstructs the trie, copying every subtrie except at the targeted key.
We fold the BinNat key into a Rebuilder—an overlay that, given an input trie, produces a new trie as its result. A Rebuilder declares the trie shape on both its input and output:
$\mathrm{Rebuilder}\mapsto\{$ $\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;$
$\mathrm{result}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\}\}$
Three observer callbacks build Rebuilders bottom-up. The $\mathrm{on\_zero}$ callback returns a Rebuilder that copies the subtries but omits the value:
$\mathrm{delete\_on\_zero}\mapsto\{$ $\mathrm{Rebuilder},\;\mathrm{result}\mapsto\{$
$\quad\mathrm{at\_odd}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd},\;$
$\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
The $\mathrm{result}.\mathrm{value}$ slot inherits only the empty declaration from $\mathrm{Rebuilder}$, effectively deleting the entry.
The $\mathrm{on\_odd}$ callback takes a Rebuilder (the recursive result for the subtrie) and returns a new Rebuilder that preserves $\mathrm{value}$ and $\mathrm{at\_even}$, but replaces $\mathrm{at\_odd}$ with the recursively rebuilt subtrie:
$\mathrm{delete\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Rebuilder},\;$
$\quad\mathrm{result}\mapsto\{\mathrm{Rebuilder},\;$
$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
$\qquad\mathrm{result}\mapsto\{$
$\qquad\quad\mathrm{value}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{value},\;$
$\qquad\quad\mathrm{at\_odd}\mapsto\mathrm{applied}.\mathrm{result},\;$
$\qquad\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
The $\mathrm{on\_even}$ callback is symmetric.
The $\mathrm{Delete}$ combinator folds a BinNat key with these callbacks to obtain a Rebuilder, then applies it to the trie:
$\mathrm{Delete}$ $\mapsto\{$
$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
$\quad\mathrm{trie\_input}\mapsto\{\},$
$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{delete\_on\_zero},\;$
$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{delete\_on\_odd},\;$
$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{delete\_on\_even}\},$
$\quad\mathrm{applied\_rebuilder}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
$\quad\mathrm{result}\mapsto\mathrm{applied\_rebuilder}.\mathrm{result}\}$
Since inheritance is idempotent, the empty scaffolding left by the Rebuilder schema merges harmlessly with any subsequent insertions.
Update
Updating the entry at key $k$ to a new value $v$ combines deletion and insertion. Because $\mathrm{Delete}$ is a combinator whose output lives in the result property, one must first project out the rebuilt trie before inheriting the new singleton. In ANF style:
$\mathrm{deleted}$ $\mapsto\{\mathrm{Delete},\;\mathrm{key}\mapsto k,\;\mathrm{trie\_input}\mapsto\mathrm{trie}\}$
$\mathrm{result}$ $\mapsto\{\mathrm{deleted}.\mathrm{result},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
where the second line uses key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ as a concrete example. The projection $\mathrm{result}$ is essential: $\mathrm{Delete}$ carries internal scaffolding ($\mathrm{key}$, $\mathrm{trie\_input}$, $\mathrm{applied\_key}$, etc.), and only its $\mathrm{result}$ property holds the rebuilt trie. This intermediate projection prevents Update from being a single symmetric inheritance.
Appendix B Church Encoding of Standard Types
This appendix demonstrates that standard data types can be encoded in Overlay-Calculus without scalar types, using Church encoding. These encodings serve as a proof of concept that a standard library can be built entirely within Overlay-Calculus. Each encoding is presented as a module overlay. In practical implementations, more efficient representations (e.g., binary naturals) or FFI to host language primitives would typically be used. The implementation (Yang, 2025b) includes executable implementations of these encodings with test cases for $\mathrm{Not}(\mathrm{True})$, $\mathrm{And}(\mathrm{True},\mathrm{True})$, $\mathrm{Or}(\mathrm{False},\mathrm{True})$, and natural number addition, all verified against expected output snapshots.
B.1. Boolean Module
The boolean module contains a schema ($\mathrm{Boolean}$), constructors ($\mathrm{True}$, $\mathrm{False}$), and operations ($\mathrm{Not}$, $\mathrm{And}$, $\mathrm{Or}$). All definitions are siblings within a single module overlay.
$\{\quad\mathrm{boolean}\mapsto\{\quad\mathrm{Boolean}$ $\mapsto\{\mathrm{on\_true}\mapsto\{\},\;\mathrm{on\_false}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
$\mathrm{True}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{True}.\mathbf{this}.\mathrm{on\_true}\},$
$\mathrm{False}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{False}.\mathbf{this}.\mathrm{on\_false}\},$
$\mathrm{Not}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{operand}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_operand}\mapsto\{\mathrm{operand},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_false},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_operand}.\mathrm{result}\},$
$\mathrm{And}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{right},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{False}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\},$
$\mathrm{Or}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{True},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{right}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\}\quad\}\quad\}$
Reference explanation
Within the $\mathrm{boolean}$ module, own properties such as $\mathrm{Boolean}$, $\mathrm{True}$, $\mathrm{False}$, $\mathrm{Not}$, $\mathrm{And}$, and $\mathrm{Or}$ are accessed via lexical references (e.g., the inheritance source $\mathrm{Boolean}$ in each definition). Qualified this is required in two situations:
  • Inherited properties. Inside True’s definition, $\mathrm{True}.\mathbf{this}.\mathrm{on\_true}$ requires qualified this because $\mathrm{on\_true}$ is inherited from Boolean via inheritance, not defined as an own property. Similarly, $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ and $\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}$ access inherited callbacks.
  • Variable shadowing. Inside Not’s $\mathrm{applied\_operand}$, the labels $\mathrm{on\_true}$ and $\mathrm{on\_false}$ are own properties of $\mathrm{applied\_operand}$ (with swapped values), so a lexical reference $\mathrm{on\_false}$ would resolve to the local definition rather than Not’s inherited callback. Qualified this $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ bypasses the shadowing.
Worked example
Consider the expression $\{\mathrm{boolean}.\mathrm{Not},\;\mathrm{operand}\mapsto\mathrm{boolean}.\mathrm{True},\;\mathrm{on\_true}\mapsto A,\;\mathrm{on\_false}\mapsto B\}$ at the root scope level, where $A$ and $B$ are arbitrary expressions. The result is an object whose $\mathrm{result}$ label contains the same observable tree as $B$, because $\mathrm{Not}$ swaps $\mathrm{on\_true}$ and $\mathrm{on\_false}$, and $\mathrm{True}$ selects $\mathrm{on\_true}$, which after swapping becomes $B$.
B.2. Natural Number Module
The natural number module contains a schema ($\mathrm{Nat}$), constructors ($\mathrm{Zero}$, $\mathrm{Succ}$), and an operation ($\mathrm{Add}$).
$\{\quad\mathrm{natural}\mapsto\{\quad\mathrm{Nat}$ $\mapsto\{\mathrm{successor}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},\;\mathrm{zero}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
$\mathrm{Zero}$ $\mapsto\{\mathrm{Nat},\;\mathrm{result}\mapsto\mathrm{Zero}.\mathbf{this}.\mathrm{zero}\},$
$\mathrm{Succ}$ $\mapsto\{\mathrm{Nat},$
$\qquad\mathrm{predecessor}\mapsto\mathrm{Nat},$
$\qquad\mathrm{applied\_predecessor}\mapsto\{\mathrm{predecessor},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{zero}\},$
$\qquad\mathrm{applied\_successor}\mapsto\{\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{argument}\mapsto\mathrm{applied\_predecessor}.\mathrm{result}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_successor}.\mathrm{result}\},$
$\mathrm{Add}$ $\mapsto\{\mathrm{Nat},$
$\qquad\mathrm{augend}\mapsto\mathrm{Nat},$
$\qquad\mathrm{addend}\mapsto\mathrm{Nat},$
$\qquad\mathrm{applied\_addend}\mapsto\{\mathrm{addend},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{zero}\},$
$\qquad\mathrm{applied\_augend}\mapsto\{\mathrm{augend},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{applied\_addend}.\mathrm{result}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_augend}.\mathrm{result}\}\quad\}\quad\}$
Concrete example
At the root scope level,
$\mathrm{one}\mapsto\{\mathrm{natural}.\mathrm{Succ},\;\mathrm{predecessor}\mapsto\mathrm{natural}.\mathrm{Zero}\}.$
Applying this with a Peano-style successor and zero: $\{\mathrm{one},\;\mathrm{successor}\mapsto S,\;\mathrm{zero}\mapsto Z\}$ produces an object whose $\mathrm{result}$ label contains the same observable tree as $\{S,\;\mathrm{argument}\mapsto Z.\mathrm{result}\}.\mathrm{result}$, i.e., successor applied once to zero. Here $S$ and $Z$ are arbitrary expressions representing a successor operation and a zero value respectively.
Appendix C Multi-Path Self-Reference in Scala
This appendix demonstrates that Scala 3 rejects the multi-path self-reference pattern that Overlay-Calculus handles naturally. Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:

### v2

@@ 8. Future Work
The Overlay language (Yang, 2025b) is an executable implementation of overlay-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
@@
The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how overlay-calculus relates to the $\lambda$-calculus: a Typed overlay-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve. In untyped overlay-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
@@
Overlay-calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5). However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
@@
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields. This does not restrict inheritance itself, which remains unconstrained, but provides a way to detect invariant violations before they propagate.
@@
The same technique, an observer returning Boolean followed by Boolean dispatch, yields closed pattern matching. The observer pattern of Section 5 is inherently open: new constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute, as they have a fixed order, and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
9. Conclusion
The goals of declarative programming forced the semantics of overlay-calculus upon us. The emergent phenomena exhibit an expressive asymmetry between overlay-calculus and the $\lambda$-calculus. Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
Appendix A Expressive Asymmetry: Proofs
Definition A.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
A construct $C$ of language $\mathscr{L}_{1}$ is macro-expressible in language $\mathscr{L}_{0}$ if there exists a translation $\mathcal{E}$ such that for every program $P$ containing occurrences of $C$, the translation $\mathcal{E}(P)$ is obtained by replacing each occurrence of $C$ with an $\mathscr{L}_{0}$ expression that depends only on $C$’s subexpressions, leaving all other constructs of $P$ unchanged.
Theorem A.2 (Forward macro-expressibility).
The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into overlay-calculus.
Proof.
Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an overlay-calculus expression that depends only on the construct’s immediate subexpressions. No rule inspects the surrounding context. ∎
Theorem A.3 (Non-expressibility of overlay).
There is no macro-expressible translation from overlay-calculus to the $\lambda$-calculus.
Proof.
Let $\mathscr{L}_{1}$ be the overlay-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of overlay trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$. Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$. Define $\cong_{1}$ on $\mathscr{L}_{1}$ as overlay-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every overlay context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
Define Church booleans via the translation $\mathcal{T}$ (Section 4): $\mathrm{true}=\lambda t.\,\lambda f.\,t$ and $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In de Bruijn notation:
$\mathrm{true}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}$
$\mathrm{false}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!0}.\mathrm{argument}\}\}$
Define Church boolean equality $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. In ANF:
$\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=b\;\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=x_{1}\;\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=a\;b\;\mathbf{in}\;x_{3}\;x_{2}$
Applying $\mathcal{T}$ (Section 4):
$\mathrm{eq}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},$
$\quad x_{1}\mapsto\{\uparrow^{\!0}.\mathrm{argument},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\quad x_{2}\mapsto\{x_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\quad x_{3}\mapsto\{\uparrow^{\!1}.\mathrm{argument},\;\mathrm{argument}\mapsto\uparrow^{\!0}.\mathrm{argument}\},$
$\quad\_\mathrm{tail}\mapsto\{x_{3}.\mathrm{result},\;\mathrm{argument}\mapsto x_{2}.\mathrm{result}\},$
$\quad\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}\}$
Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
In overlay-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$. Define overlay trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
$e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
$e_{2}\;=\;\{$ $\mathrm{eqTrue}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqTrue}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
That is, $e_{1}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{false}\;\mathrm{false}))$ and $e_{2}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{true}\;\mathrm{true}))$, where $\mathrm{ANF}$ converts a $\lambda$-term to A-normal form by naming every intermediate application (Section 4). Since both $\mathrm{ANF}$ and $\mathcal{T}$ are macro-expressible, their composition is too, so $e_{1},e_{2}\in\mathscr{L}_{0}$. Both $e_{1}.\mathrm{result}$ and $e_{2}.\mathrm{result}$ evaluate to Church $\mathrm{true}$. By Full Abstraction (Theorem 4.3), $\mathrm{BT}(e_{1})=\mathrm{BT}(e_{2})$, so $e_{1}\cong_{0}e_{2}$.
@@ Appendix A Expressive Asymmetry: Proofs
$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope. Thus $\mathrm{firstOperand}$ projects $\lambda a$’s argument and $\mathrm{secondOperand}$ projects $\lambda b$’s argument.
In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope. By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings. Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$. These produce structurally distinct overlay trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the overlay facilities of $\mathscr{L}_{1}$. ∎
Corollary A.4.
Overlay-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
Appendix B Multi-Path Self-Reference in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that overlay-calculus handles naturally.
Scala 3
Consider two objects that independently extend an outer class, each providing its own copy of an inner trait: